Find the linear substitution that reduces the inhomogeneous equation to the homogeneous one: y,=x+y-3/y-x+1


1
Expert's answer
2021-11-29T17:07:24-0500

A first order differential equation

"dy\/dx=f(x,y)"

is called homogeneous equation, if the right side satisfies the condition

"f(tx,ty)=f(x,y)"


we have:

"f(x,y)=x+\frac{y-3}{y-x}+1"

then:

"f(tx,ty)=tx+\frac{ty-3}{ty-tx}+1"


"x+\frac{y-3}{y-x}+1=tx+\frac{ty-3}{ty-tx}+1"


"x(y-x)+y-3=t^2x(y-x)+ty-3"

"x(y-x)(t^2-1)+y(t-1)=0"

"x(y-x)(t+1)+y=0"


"t=-\frac{y}{x(y-x)}-1"